# Implementing a Robust Test of Galaxy Catalogue Completeness for Dark Siren Measurements of the Hubble Constant

### v1

@@ Abstract
We present the application of a robust test of galaxy catalogue completeness to the gwcosmo pipeline. The method implements a straightforward statistical test for determining the apparent magnitude completeness limit of a magnitude-redshift sample. This offers an improved, less conservative approach compared with how galaxy catalogue completeness is currently estimated in the gwcosmo gravitational wave cosmology pipeline for determining the Hubble constant $H_{0}$. The test also does not require prior knowledge of the luminosity function, and thus returns a more robust estimate of the limiting apparent magnitude for a magnitude-redshift sample of galaxies. For GWTC-1 results using $B_{J}$-band photometry of galaxies in the GLADE catalogue, we find a $3.4\%$ improvement on the inference of $H_{0}$ using dark sirens only and a $1.3\%$ improvement for the combined posterior with GW170817. Using GLADE+, there is a $8.6\%$ improvement with dark sirens only and a $6.3\%$ improvement for the combined posterior with GW170817. However, the final posterior on $H_{0}$ using the GWTC-3 dataset with the GLADE+ $K$-band shows no improvement when applying the robust method. This is because the GLADE+ galaxy catalogue provides little or no coverage in the $K$-band for any of the GWTC-3 events. With the use of deeper galaxy catalogues in future gravitational wave cosmology analyses, the adoption of a less conservative estimate of magnitude completeness will become increasingly important.
@@ 3.1 GWTC-1
The GWTC-1 data has previously been used in the inference of $H_{0}$ using the GLADE catalogue. Here we repeat the analysis using the $B_{J}$-band of both the GLADE and GLADE+ catalogues. One caveat to note is that the assumption that the luminosity function is universal might fail when using the $B_{J}$-band of the GLADE+ catalogue; therefore, this analysis should serve as proof of principle for consequent analyses.
@@ 3.2 GWTC-3
The GWTC-3 catalogue consists of 90 events, of which 47 are used in this analysis. Of these 47 events, 42 are BBHs, 2 are BNSs (the bright siren GW170817 and GW190425), 2 are NSBHs (GW200105 and GW200115) and one is the asymmetric mass binary GW190814 (Abbott et al., 2023a, 2017a, 2020a, 2021b, 2020b). In the original analysis presented in Abbott et al. (2023b), the $K$-band of the GLADE+ catalogue was found to be more appropriate than the $B_{J}$-band for analysis; it is less affected by galactic dust, and the behaviour of its luminosity function can therefore be better approximated. While this does not affect tests of galaxy catalogue completeness, it does affect the luminosity weighting of galaxies in the sample (Abbott et al., 2023b).
The completeness of the GLADE+ catalogue decreases more rapidly past $d_{L}\sim 100$Mpc in the $K$-band than in the $B$-band (Dálya et al., 2018). Applying the robust method to the GLADE+ catalogue with a pixel size of $N_{\rm side}=32$, the mean $K$-band apparent magnitude threshold is $m_{\rm thr}=13.49$. By comparison, the median method gives $m_{\rm thr}=12.91$. While the robust method allows us to use more galaxies, the apparent magnitude threshold is still comparatively bright, reducing the impact of the method on the recovered $H_{0}$ compared to when using the $B_{J}$-band.
@@
Figure 2: Final posterior on $H_{0}$ using the GLADE $B_{J}$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
Figure 3: Final posterior on $H_{0}$ using the GLADE+ $B_{J}$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
@@
Figure 5: Results using the robust (blue) and median (orange) methods for individual events using the GLADE+ $B_{J}$-band with the GWTC-1 catalogue. The event GW170814 is analysed using the $g$-band of the DES-Y1 catalogue.
@@
Table 1: Final results for constraints on $H_{0}$ in km s-1 Mpc-1 from the GWTC-1 and GWTC-3 datasets, using different methods and galaxy catalogues. The GWTC-1 dataset was analysed using the $B_{J}$-band of both the GLADE and GLADE+ catalogues, while the GWTC-3 dataset was analysed using the $K$-band of the GLADE+ catalogues. Confidence intervals are quoted at the $1\sigma$ level.
@@ 4 Conclusions
However, the final posterior on $H_{0}$ showed significant improvement when applying the robust method to the GWTC-1 analysis using the $B_{J}$-band of the GLADE+ catalogue. When only dark sirens were considered, the $1\sigma$ posterior was 8.6% narrower when using the robust method than when using the median apparent magnitude as a threshold. While the GLADE+ $B$-band is less reliable for tests of completeness due to the behaviour of its luminosity function [^2] See the discussion in (Abbott et al., 2023b), the result demonstrates the need for a careful treatment of the apparent magnitude threshold of future, deeper galaxy catalogues in order to obtain the best constraints on $H_{0}$ from dark standard sirens.
@@
Our future work will focus on applying the robust method to mock data in order to fully characterise potential biases and explore the effect of a more rigorously and robustly defined $m_{\rm thr}$ on the inference of $H_{0}$ when analysing GW data with deeper galaxy catalogues. We will also extend analysis to other colour bands, including the $B_{J}$ band, seeking to exploit the property of the robust method that it does not require the adoption of a specific parametric form for the galaxy luminosity function. Moreover, our future work we will also extend the analysis presented here to the case of galaxy surveys described by both a faint and bright apparent magnitude limit, applying the robust completeness test first developed in Johnston et al. (2007).

### v2

@@ Abstract
We present the application of a robust test of galaxy catalogue completeness to the gwcosmo pipeline. The method implements a straightforward statistical test for determining the apparent magnitude completeness limit of a magnitude-redshift sample. This offers an improved, less conservative approach compared with how galaxy catalogue completeness is currently estimated in the gwcosmo gravitational wave cosmology pipeline for determining the Hubble constant $H_{0}$. The test also does not require prior knowledge of the luminosity function, and thus returns a more robust estimate of the limiting apparent magnitude for a magnitude-redshift sample of galaxies. For GWTC-1 results using $B$-band photometry of galaxies in the GLADE catalogue, we find a $3.4\%$ improvement on the inference of $H_{0}$ using dark sirens only and a $1.3\%$ improvement for the combined posterior with GW170817. Using GLADE+, there is a $8.6\%$ improvement with dark sirens only and a $6.3\%$ improvement for the combined posterior with GW170817. However, the final posterior on $H_{0}$ using the GWTC-3 dataset with the GLADE+ $K$-band shows no improvement when applying the robust method. This is because the GLADE+ galaxy catalogue provides little or no coverage in the $K$-band for any of the GWTC-3 events. With the use of deeper galaxy catalogues in future gravitational wave cosmology analyses, the adoption of a less conservative estimate of magnitude completeness will become increasingly important.
@@ 3.1 GWTC-1
The GWTC-1 data has previously been used in the inference of $H_{0}$ using the GLADE catalogue. Here we repeat the analysis using the $B$-band of both the GLADE and GLADE+ catalogues. One caveat to note is that the assumption that the luminosity function is universal might fail when using the $B$-band of the GLADE+ catalogue; therefore, this analysis should serve as proof of principle for consequent analyses.
@@ 3.2 GWTC-3
The GWTC-3 catalogue consists of 90 events, of which 47 are used in this analysis. Of these 47 events, 42 are BBHs, 2 are BNSs (the bright siren GW170817 and GW190425), 2 are NSBHs (GW200105 and GW200115) and one is the asymmetric mass binary GW190814 (Abbott et al., 2023a, 2017a, 2020a, 2021b, 2020b). In the original analysis presented in Abbott et al. (2023b), the $K$-band of the GLADE+ catalogue was found to be more appropriate than the $B$-band for analysis; it is less affected by galactic dust, and the behaviour of its luminosity function can therefore be better approximated. While this does not affect tests of galaxy catalogue completeness, it does affect the luminosity weighting of galaxies in the sample (Abbott et al., 2023b).
The completeness of the GLADE+ catalogue decreases more rapidly past $d_{L}\sim 100$Mpc in the $K$-band than in the $B$-band (Dálya et al., 2018). Applying the robust method to the GLADE+ catalogue with a pixel size of $N_{\rm side}=32$, the mean $K$-band apparent magnitude threshold is $m_{\rm thr}=13.49$. By comparison, the median method gives $m_{\rm thr}=12.91$. While the robust method allows us to use more galaxies, the apparent magnitude threshold is still comparatively bright, reducing the impact of the method on the recovered $H_{0}$ compared to when using the $B$-band.
@@
Figure 2: Final posterior on $H_{0}$ using the GLADE $B$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
Figure 3: Final posterior on $H_{0}$ using the GLADE+ $B$-band on the GWTC-1 dataset. Vertical dashed lines show the $1\sigma$ intervals. The solid lines show results showing the robust method, while dash-dotted lines show results using $m_{\rm med}$ as the apparent magnitude threshold.
@@
Figure 5: Results using the robust (blue) and median (orange) methods for individual events using the GLADE+ $B$-band with the GWTC-1 catalogue. The event GW170814 is analysed using the $g$-band of the DES-Y1 catalogue.
@@
Table 1: Final results for constraints on $H_{0}$ in km s-1 Mpc-1 from the GWTC-1 and GWTC-3 datasets, using different methods and galaxy catalogues. The GWTC-1 dataset was analysed using the $B$-band of both the GLADE and GLADE+ catalogues, while the GWTC-3 dataset was analysed using the $K$-band of the GLADE+ catalogues. Confidence intervals are quoted at the $1\sigma$ level.
@@ 4 Conclusions
However, the final posterior on $H_{0}$ showed significant improvement when applying the robust method to the GWTC-1 analysis using the $B$-band of the GLADE+ catalogue. When only dark sirens were considered, the $1\sigma$ posterior was 8.6% narrower when using the robust method than when using the median apparent magnitude as a threshold. While the GLADE+ $B$-band is less reliable for tests of completeness due to the behaviour of its luminosity function [^2] See the discussion in (Abbott et al., 2023b), the result demonstrates the need for a careful treatment of the apparent magnitude threshold of future, deeper galaxy catalogues in order to obtain the best constraints on $H_{0}$ from dark standard sirens.
@@
Our future work will focus on applying the robust method to mock data in order to fully characterise potential biases and explore the effect of a more rigorously and robustly defined $m_{\rm thr}$ on the inference of $H_{0}$ when analysing GW data with deeper galaxy catalogues. We will also extend analysis to other colour bands, including the $B$ band, seeking to exploit the property of the robust method that it does not require the adoption of a specific parametric form for the galaxy luminosity function. Moreover, our future work we will also extend the analysis presented here to the case of galaxy surveys described by both a faint and bright apparent magnitude limit, applying the robust completeness test first developed in Johnston et al. (2007).